OpenStudy (anonymous):

Find the center and radius of the circle whose equation is x^2+ y^2-20 4x- 2y .

OpenStudy (anonymous):

is it x^2 -4x + y^2 +2y-20 = 0 (x-2)^2 (y+1)^2 -20 = 0 + 2 + 1 x^2-4x+4 y^2-2y+1-20 =3 x^2-4x + y^2 -2y +5-20 x^2-4x +y^2-2y-15=3 x(x-4) y(y-2)-15=3 is this the right direction so far?

OpenStudy (campbell_st):

you need to complete the square for terms in x and y the standard form is (x -h)^2 + (y = k)^2 = r^2 the centre of the circle is (h, k) and the radius is r so you have x^2 - 4x + (4/2)^2 + y^2 + 2y + (2/1)^2 = 20 + (4/2)^2 + (2/1)^2 they terms in x and y can be factorised into perfect squares and the right hand side will give r^2 when simplified.

OpenStudy (anonymous):

so it should be (x+2)^2 + (y+1)^2 = 23

OpenStudy (campbell_st):

not quite the right hand side should be 20 + 4 + 1 the term in x has a negative middle term so the sign inside the perfect square should be negative...

OpenStudy (anonymous):

im srry my brain is crashing been studying for a while. I see the mistake with the negative and kinda with the 20+4+1 but my completing the square is rusty isnt it taking 1/2 of b and adding it on both sides then bringing it down to rewrite it so when its squared it gives u the statement above?

OpenStudy (campbell_st):

ok... so you have (x - 2)^2 + (y + 1)^2= 25 use the information above to find the centre and radius

OpenStudy (anonymous):

well the radius is 5 and the center is 2,-1 but Im still not seeing how u got 25 lol

OpenStudy (anonymous):

I keep getting 23

OpenStudy (anonymous):

cause half of b for the y part should be 1

OpenStudy (campbell_st):

ok... complete the square for x^2 - 4x b = -4 the constant term in the perfect square is (-4/2)^2 so you are really doing (-2)^2 = 4 add it to both sides of the equation to keep it in balance... same for y^2 + 2y b = 2 (2/2)^2 = 1 so add 1 to both sides to keep the equation in balance so you have x^2 -4x + 4 + y^2 +2y + 1 = 20 + 4 + 1 (x -2)^2 + (y + 1)^2 = 25 hope that makes sense